40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 3 22:41:17 2019, shruti, Update, General, PLL for PM measurement  PM_measurement.jpeg
    Reply  Wed Dec 4 20:32:43 2019, gautam, Update, General, PLL for PM measurement  
    Reply  Mon Dec 9 21:22:46 2019, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 
       Reply  Tue Dec 17 20:08:09 2019, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 20191217.png
          Reply  Thu Jan 16 19:32:23 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT 
             Reply  Thu Jan 23 20:08:49 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT Image-1.jpg
                Reply  Tue Jan 28 19:40:15 2020, shruti, Update, General, PLL / PM measurement of Xend NPRO PZT PMTF.pdfBeatSpectrum.pdf
Message ID: 15129     Entry time: Thu Jan 16 19:32:23 2020     In reply to: 15101     Reply to this: 15148
Author: shruti 
Type: Update 
Category: General 
Subject: PLL / PM measurement of Xend NPRO PZT 

With Gautam's help today the PLL managed to be be locked for a few brief moments. Turns out the signal power of the beat was an issue.

What was changed prior to/ during the experiment:

1. The PSL shutter was closed so not light goes into the input mode cleaner.

2. HEPA turned up (will be turned back down to ~30%)

3. AOM driver offset voltage decreased from 1V to ~100 mV (this will be reverted to 1V by the end of today). This increases the beat signal by deflecting the zeroth order beam to create the beat.

4. Output of servo SR 560 sent to the PZT of the X NPRO laser (the cable was disconnected from the pomona box at the X end)

5. The SR560, mixer, LPF and cables required for connections were moved into the PSL enclosure.

6. The error and control signals were hooked up to the oscilloscope where the beat outputs were visible (the setup has been reverted back to the original).

 

Elog 14687 has a detailed description of the conditions that provide a stable lock. I was told that the PI controller (LB1005) may be a better servo than the SR560, but today it was not used.

1) Parameters during the more successful attempts:

LPF: 5 MHz, Mixer: ZP-3+

Gain set at SR560: varied, but generally 200

Filter at SR560: 1 Hz low pass (single pole? at least by the label)

2) The LO had to be very close (<2 MHz) to the beat frequency in order to achieve a lock for ~30s


gautam edits:

  • the error signal for the PLL was being sourced from the 20dB coupled port on the BeatMouth.
  • additionally, most of the power in the PSL beam coupled into the fiber was being deflected into the first order beam by team ringdown.
  • The Vpp of the mixer output (when using the coupled beat and low PSL beam power) was a paltry 5-10 mVpp no.
  • I suggested using the direct NF1611 output for this measurement instead of the coupled output (alternatively, use an amp). it's probably also better to use the LB1005 for locking the PLL, long term, this can be set up to be controlled remotely, and a slow PID servo can be used to extend the duration of the lock by servoing either the marconi carrier freq or the EX temp ctrl.
Quote:

1. Some calculations

For a Unity Gain Frequency (UGF) of 1 kHz, assumed PZT response K_{VCO} of 1 MHz/V, Mixer response K_{M} of 25 mV/\pi rad, the required gain of the amplifier is

G = 2 \pi \times \text{UGF}/ (K_{VCO} K_M)

G ~ 0.8

2. Progress

- Measured the mixer response

Measuring mixer response:

- PSL laser temperature was adjusted so that beat frequency was roughly 25 MHz and the amplitude was found to be roughly -30dBm.

- At the RF port instead of the beat signal, a signal of 25 MHz + few kHz at -30 dBm was inputted. The LO was a 25 MHz signal was sent from the Marconi at 7 dBm.

- The mixer output was measured, with setup as in Attachment 1  Figure (A), on an oscilloscope. The slope near the small angle region of the sine curve would be the gain (in V/rad) and was found to be: K_M \approx 25 \text{ mV}/ \pi rad

- Since from the above calculations it seemed like an amplifer gain of 1 should work for the PLL, I rearranged the set up as in Figure (B) of Attachment 1 to actuate the X end NPRO PZT, I adjusted the PSL temperature (slow control) to try and match the frequency to 25 MHz, but couldn't lock the loop. I was monitoring the error signal after amplification (50 ohm output of the SR 560) which showed oscillations when the beat frequency was near 25 MHz and nothing significant otherwise.

- I used a 20 dB attenuator at the amplifier output and saw the beat note oscillate for longer, but maybe because it was a 50 ohm component in a high impedance channel it did not work either (?). I tried other attenuator combinations with no better luck.

- Is there a better location to add the attenuator? Should I pursue amplifying the beat signal instead?

- Also, it seemed like the beat note drift was higher than earlier. Could it be because the PMC was unlocke

ELOG V3.1.3-